# Fundamental physical constants set the observability and operation of phase and other transitions and increase entropy

### v2

@@ Abstract
Approaching the problem of understanding fundamental physical constants (FPCs) started with discussing the role these constants play in high-energy nuclear physics and astrophysics. Condensed matter physics was relatively unexplored in this regard. More recently, it was realised that FPCs set lower or upper bounds on key condensed matter properties. In this Perspective, we discuss a wider role played by FPCs in condensed matter physics: at given environmental conditions, FPCs set the observability and operation of entire physical effects and phenomena. We discuss first and second-order phase transitions and emerging new states including long-lived metastable states. We also discuss transitions between metastable states, chemical reactions and their products. These examples show that different values of FPCs would disable these transitions, and new phases as a result of these transitions would not emerge. This suggests that the current values of FPCs, by enabling transitions which give rise to new states, act to increase the phase space and entropy.
@@ 1 Introduction
Fundamental physical constants (FPCs), determine physical processes at all energy and length scales of our experiments and theories. These different length scales are illustrated in Figure 1. The values of FPCs are often listed in textbooks and reviews (see, e.g., Refs. Uzan (2003, 2011); Tiesinga et al. (2021); Ashcroft and Mermin (1976)), and their recommended values are updated and maintained in the National Institute for Standards and Technology database NIST .
@@
The top in Figure 1 involves cosmological length scales and a balance between the gravitational constant and the cosmological constant. This balance gives enough time for the universe to expand and not to contract too quickly and at the same time maintain sufficient gravitational clamping so that galaxies, stars and other structures can form Barrow (2003); Adams (2019).
@@
In this Perspective, we discuss a wider role of FPCs in condensed matter: extending beyond setting bounds on properties that are already observed, fundamental constants affect observability and operation of entire physical phenomena and effects. In the next section, we discuss why discussing FPCs at the condensed matter scale is important. This is followed by several examples in condensed matter physics showing how FPCs set observability and operation of phase transitions, the existence of different phases and states of matter emerging as a result of these transitions, including numerous metastable states. We also discuss chemical reactions and their products. These examples show that were FPCs to take different values, transitions would be disabled at our environmental conditions and new different phases would not emerge. This suggests that the current values of FPCs, by enabling transitions and new states emerging as a result of these transitions, act to increase the phase space and entropy. This invites an analogy with an evolutionary process where entropy increase selects for FPCs.
@@ 2 Why filling the condensed matter gap is important
3 Observability and operation of phase and other transitions. Existence of different phases and states
@@ 3.1 Fundamental and environmental parameters
We start with separating two types of parameters. The first is the set of FPCs. The second type are the environmental parameters such as, for example, temperature and pressure at typical planetary conditions where we do our observations. These parameters are not fundamental but are due to historical accidents Weinberg (1993) such as the distance between the planet and its star which affects the temperature of the planet. In several examples in this section, we consider a certain range of environmental parameters (for example, the temperature range 102-103 K at typical planetary conditions in the Solar system) and ask how the observability and operation of condensed matter phenomena change as a result of different FPCs.
@@
$\alpha$ and $\beta$ are of particular importance when discussing processes at different length scales in Figure 1, including the stability of particles, nuclear reactions, formation and evolution of stars, synthesis of heavy nuclei and so on. These processes impose constraints on the allowed range of $\alpha$ and $\beta$ (as well as on the strong nuclear constant) Barrow (2003); Adams (2019).
It is useful at this stage to recall that characteristic scales of distance and energy in condensed matter states, solids and liquids, are set by the Bohr radius $a_{\rm B}$ and Rydberg energy $E_{\rm R}$ as
@@
$a_{\rm B}$ is about 0.5 Å and gives a shorter range of typical interatomic separations condensed phases in the range 1-3 Å. $E_{\rm R}$ is about 13.6 eV and gives a larger value of cohesive energy in condensed phases in the typical range 1-10 eV.
@@ 3.2 Second-order phase transitions
Our first case study are the second-order phase structural transitions in solids. We consider common silica SiO2 and its low-temperature phase, quartz. Silica has several polymorphs at different pressure and temperature, separated by second-order phase transitions. Let’s first consider $\beta$-cristobalite, the high-temperature phase emerging at the phase transition around 1800 K Hammonds et al. (1996); Dove (1997).
@@
Using $\langle\Delta p_{i}\rangle^{2}\propto\frac{T}{\omega_{i}^{2}}$ and $\langle\Delta q_{i}\rangle^{2}\propto\frac{T}{\omega_{i}^{2}}$ for each phonon mode in Eq. (6) and taking the relation between $\Delta\Gamma$ and $S$ as
@@
Cristobalite is a long-lived metastable phase of silica which can be cooled back to room temperature or any other temperature below the phase transition Hammonds et al. (1996); Dove (1997). According to Eq. (5), there is no net entropy change due to heating quartz from temperature $T_{1}$ to the transition temperature $T_{2}$ and cooling back the metastable cristobalite phase: the entropy first increases by $3N\ln\frac{T_{2}}{T_{1}}$ and then decreases by the same amount (we do not consider the small effect of phonon frequency softening with temperature due to usual anharmonicity). Before proceeding further, we add a note about the entropy and other thermodynamic functions of non-equilibrium metastable systems. These functions are defined in the commonly used sense: at times $\Delta t$ shorter than the relaxation time of the non-equilibrium system, this system can be conditionally considered in an equilibrium state where observed properties agree with those described by equilibrium thermodynamic functions Landau and Lifshitz (1970). This is discussed in Section 3.5 in more detail.
As a more general case compared to pure metastable cristobalite, let’s consider the entropy of the quartz-cristobalite mixture at room or any other temperature below the transition point where both phases co-exist, $S_{Q-C}$. This mixture consists of any proportion of the two phases with the total number of atoms $N$. We compare $S_{Q-C}$ to the entropy of the initial quartz phase, $S_{Q}$, with the same number of atoms $N$. $S_{Q-C}$ is larger than $S_{Q}$ because the entropy of the cristobalite phase is larger as discussed earlier. Hence,
$S_{Q-C}>S_{Q}$ (8)
We therefore find that the entropy increase (8) is ultimately related to a phase transition which gives rise to the second higher-entropy phase. The increased phase space and entropy in the high-temperature phase become “frozen-in” in the metastable state at low $T$.
@@
Without a phase transition operating in the considered range, the entropy would increase according to Eq. (5) within the single quartz phase but without the higher entropy contribution from the second metastable phase.
@@
We now observe that if we consider different values of FPCs, the phase transition would not operate. The upper bound on the maximal phonon frequency $\omega_{m}$ in condensed matter can be written in terms of fundamental constants Trachenko et al. . This bound follows from combining Eq. (4) and the ratio of $\hbar\omega_{m}$ and cohesive energy $E$ as
@@
where $m$ is the atom mass.
Setting $E=E_{\rm R}$ gives
@@
which has the upper bound, $\omega_{m}^{u}$, for $m=m_{p}$ as Trachenko et al.
@@
Lets consider what would happen to a phase transition were FPCs to take different values and in particular consider variations of FPCs which increase $\omega_{m}$ and $\omega_{m}^{u}$ in Eqs. (10)-(11). Note that this question is well-posed because it is possible to change $\omega_{m}$ and $\omega_{m}^{u}$ by varying $\hbar$, $m_{e}$, $e$ and $m_{p}$ while keeping $\beta$ and $\alpha$ in Eqs. (1) and (2) unchanged and hence maintain all preceding essential processes (e.g. those involved in stellar formation and evolution, synthesis of heavy elements, stability of ordered structures and so on) intact Barrow (2003); Adams (2019).
Larger $\omega_{m}$ implies larger $\omega_{i}$ and $\omega_{\rm D}$ (recall that $\omega_{m}$ is close to $\omega_{\rm D}$) and $T\ll\hbar\omega_{\rm D}$ in the temperature range considered. This, in turn, implies that the high-temperature classical result (5) is no longer applicable because the system is now in the quantum regime. In this regime, instead of Eq. (5), the entropy is Landau and Lifshitz (1970)
@@
$S_{s}$ is close to 0 when $T\ll\hbar\omega_{\rm D}$. This means that the second term in the free energy $F=E-TS$ is close to 0 too, and $F$ now does not have a crossover corresponding to a phase transition. The phase transition becomes inoperative.
Therefore, the current values of fundamental constants, by enabling the operation of phase transitions which give rise to new states, increase the phase space and entropy in Eq. (8).
A binary quartz-cristobalite mixture with increased entropy (8) is readily generalised to a collection of many systems existing in long-lived metastable states. Figure 2 illustrates this point. The increase of the phase space and entropy in these metastable states is generally expected because (a) high-temperature phases tend to be more symmetric Landau and Lifshitz (1970) with more open structure and have a larger number of soft phonons as a result Hammonds et al. (1996); Dove (1997) and (b) the increased phase space and entropy in the high-temperature phases become “frozen-in” in metastable states at low $T$.
Figure 2: Quartz on the left transforms to cristobalite at high temperature as a result of a phase transition. The mixture of quartz and metastable cristobalite at low temperature has combined phase space and entropy that are larger than those of quartz alone (with the same number of atoms). Phase transitions between different states enable sets of long-lived metastable systems with higher combined entropy at the same temperature. Shown minerals are for illustration only.
This case study gives us several insights. First, FPCs govern key effects in condensed matter physics and define the regime in which the system is, classical or quantum. Second, FPCs affect the observability and operation of entire effects such as phase transitions. Third, FPCs, by enabling phase transitions, can increase the phase space and entropy.
@@ 3.4 First-order phase transitions: solids, liquids, gases
FPCs resulting in large cohesive energies and melting temperatures imply that matter would exist in the solid state only at given environmental conditions. Melting and boiling phase transitions would be unobservable. Conversely, FPCs resulting in low cohesive energies and melting temperatures imply that matter would exist in the gas state only.
@@
Differently to solid and gas entropies we considered, no closed equation for the liquid entropy exists which corresponds to experimental properties of real liquids such as $c_{v}=3$ close to melting ($c_{v}$ is the constant-volume specific heat). This is due to the fundamental no-small parameter problem of the liquid theory (see Ref. Trachenko (2024) for a quick review and Refs. Trachenko and Brazhkin (2016); Trachenko (2023c) for a more detailed discussion). Existing models such hard-sphere and Van der Waals models and their extensions give $c_{v}=\frac{3}{2}$ and hence describe the ideal gas thermodynamically. However, the energy, the primary property in statistical physics Landau and Lifshitz (1970), can be calculated in liquids in the closed form using the main premise of statistical physics: energy of the system is the sum of energies of its excitations Lifshitz and Pitaevskii (2006). In liquids, these excitations are gapped phonons and particle jumps Trachenko and Brazhkin (2016); Trachenko (2023c); Proctor (2020, 2021); Chen (2022); de With (2024). The entropy can then be found by integrating $\frac{c_{v}}{T}$, where $c_{v}$ is calculated from the energy. Nevertheless, liquids are a condensed state of matter, and our discussion of how the interplay of cohesive energies set by FPCs and environmental parameters gives rise to different states of matter applies to liquids too.

### v3

@@ Abstract
Approaching the problem of understanding fundamental physical constants (FPCs) started with discussing the role these constants play in high-energy nuclear physics and astrophysics. Condensed matter physics was relatively unexplored in this regard. More recently, it was realised that FPCs set lower or upper bounds on key condensed matter properties. In this Perspective, we discuss a wider role played by FPCs in condensed matter physics: at given environmental conditions, FPCs set the observability and operation of entire physical effects and phenomena. We discuss second-order phase transitions including structural and superconducting transitions and first-order transitions including melting and boiling. We also discuss metastable states, transitions between them, chemical reactions and their products. An interesting byproduct of this discussion is that the order of magnitude of the transition temperature can be calculated from FPCs. We show that the new states emerging as a result of various transitions increase the phase space and entropy. Were FPCs to take different values, these transitions would become inoperative at our environmental conditions and the new states due to these transitions would not emerge. Generalising this insight to other levels of structure in the Universe, this suggests that the current values of FPCs, by enabling various transitions and reactions which give rise to new states, act to increase the phase space and entropy.
@@ 1 Introduction
Fundamental physical constants (FPCs) determine physical processes at all energy and length scales of our experiments and theories. These different length scales are illustrated in Figure 1. The values of FPCs are often listed in textbooks and reviews (see, e.g., Refs. Uzan (2003, 2011); Tiesinga et al. (2021); Ashcroft and Mermin (1976)), and their recommended values are updated and maintained in the National Institute for Standards and Technology database NIST .
@@
The top in Figure 1 involves cosmological scales and a balance between the gravitational constant and the cosmological constant. This balance gives enough time for the Universe to expand and not to contract too quickly and at the same time maintain sufficient gravitational clamping so that galaxies, stars and other structures can form Barrow (2003); Adams (2019).
@@
In this Perspective, we discuss a wider role of FPCs: extending beyond setting bounds on properties that are already observed, fundamental constants affect observability and operation of entire physical phenomena and effects. In the next section, we discuss why discussing FPCs at the condensed matter scale is important. This is followed by considering second-order phase transitions including structural and superconducting transitions and first-order transitions including melting and boiling. This discussion shows how FPCs set the observability and operation of phase transitions and hence the existence of different states emerging as a result of these transitions, including metastable states. An interesting byproduct of this discussion is that the order of magnitude of the transition temperature in, for example, structural phase transitions, can be calculated from FPCs.
We show that the new states emerging from various transitions increase the phase space and entropy. We also show that were FPCs to take different values, these transitions would become inoperative at our environmental conditions, and the new states due to these transitions would not emerge. Generalising this insight to other levels of structure in the Universe, this suggests that the current values of FPCs, by enabling various transitions and reactions which give rise to new states, act to increase the phase space and entropy. This invites an analogy with an evolutionary process where entropy increase selects for FPCs.
@@ 2 Why filling the condensed matter gap is important
3 Observability and operation of phase and other transitions. Emergence of new states and entropy
@@ 3.1 Fundamental and environmental parameters
We start with separating two types of parameters. The first type are FPCs. The second type are the environmental parameters such as temperature and pressure at typical planetary conditions where we do our observations. These parameters are not fundamental but are due to historical accidents Weinberg (1993) such as the distance between the planet and its star which affects the planet’s temperature. In several examples in this section, we consider a certain range of environmental parameters ($\Delta T$,$\Delta P$) around, for example, temperature (102-103 K) and pressure range at typical planetary conditions in the Solar system and ask how the observability and operation of condensed matter phenomena change as a result of different FPCs.
@@
$\alpha$ and $\beta$ are important in key processes at different length and energy scales in Figure 1, including the stability of particles, nuclear reactions, formation and evolution of stars, synthesis of heavy nuclei and so on. These processes impose constraints on the allowed range of $\alpha$ and $\beta$ (as well as on the strong nuclear constant) Barrow (2003); Adams (2019).
It is useful at this stage to recall that characteristic scales of distance and energy in condensed matter are set by the Bohr radius $a_{\rm B}$ and Rydberg energy $E_{\rm R}$ as
@@
$a_{\rm B}\approx 0.5$ and is on the short side of interatomic separations in condensed phases in the typical range 1-3 Å. $E_{\rm R}\approx$ 13.6 eV and is on the large side of cohesive energies in condensed phases in the range 1-10 eV.
@@ 3.2 Second-order phase transitions
3.2.1 Enropy increase due to a phase transition
Our first case study are second-order phase structural transitions in solids. Let’s consider common silica SiO2 and its low-temperature phase, quartz. Silica has several polymorphs at different pressure and temperature, separated by second-order phase transitions. As a model example, let’s the high-temperature phase of silica, $\beta$-cristobalite, emerging at the phase transition Hammonds et al. (1996); Dove (1997).
@@ 3.2.1 Enropy increase due to a phase transition
Using $\langle\Delta p_{i}\rangle^{2}\propto\frac{T}{\omega_{i}^{2}}$ and $\langle\Delta q_{i}\rangle^{2}\propto\frac{T}{\omega_{i}^{2}}$ for each phonon mode in Eq. (6) and
@@
Frequency softening above the phase transition can be seen more generally in terms of the order parameter $\eta$. Considering that phonon frequencies are renormalised by anharmonicity related to a phase transition, the renormalised frequencies $\tilde{\omega}_{i}$ are Dove (1997, 2003):
$\tilde{\omega}_{i}^{2}=\omega_{i}^{2}+\frac{1}{2}\delta_{i}\eta^{2}$ (8)
where $\omega_{i}$ are harmonic frequencies and $\delta_{i}$ are the anharmonicity coefficients.
In the high-temperature phase above the phase transition temperature $T_{c}$, $\eta=0$. Phonon frequencies become smaller as a result, increasing the entropy.
@@
Importantly, the entropy increase is ultimately related to a phase transition which gives rise to the second high-entropy phase. Without a phase transition operating, the entropy would increase according to Eq. (5) within the single low-temperature phase only but without larger entropy coming from the second high-temperature phase.
The entropy increase due to the phase transition can take place both above and below $T_{c}$. For reversible phase transition where the low-temperature phase is always recovered below $T_{c}$, the entropy increase takes place above $T_{c}$. In our model quartz-cristobalite example, this gives
$S_{C}(T>T_{c})>S_{Q}(T>T_{c})$ (9)
where $S_{C}(T>T_{c})$ is the entropy of the high-temperature cristobalite phase above $T_{c}$ and $S_{Q}(T>T_{c})$ is the entropy that quartz would have above $T_{c}$ without the phase transition operating. More generally, (9) applies to any two low and high-temperature phases.
The entropy increase can also take place below $T_{c}$ if the high-temperature phase becomes a long-lived metastable state below $T_{c}$. In our model quartz-cristobalite example, the phase transition is a reconstructive phase transition with large energy barriers separating the two phases. As a result, cristobalite is a long-lived metastable phase which can be cooled back to room temperature or any other temperature below the phase transition Hammonds et al. (1996); Dove (1997). According to Eq. (5), there is no net entropy change due to heating quartz from temperature $T_{1}$ to the transition temperature $T_{2}$ and cooling back the metastable cristobalite phase: the entropy first increases by $3N\ln\frac{T_{2}}{T_{1}}$ and then decreases by the same amount (we do not consider the small effect of phonon frequency softening with temperature due to usual anharmonicity). Before proceeding further, we add a note about the entropy and other thermodynamic functions of non-equilibrium metastable systems. These functions are defined in the commonly used sense: at times shorter than the relaxation time of the non-equilibrium system, this system can be conditionally considered in an equilibrium state where observed properties are described by equilibrium thermodynamics Landau and Lifshitz (1970). This is discussed in Section 3.5 in more detail.
We find both quartz and cristobalite at our environmental conditions, hence let’s consider the entropy of the quartz-cristobalite mixture at room or any other temperature below $T_{c}$ where both phases co-exist, $S_{Q-C}$. This mixture consists of any proportion of the two phases with the total number of atoms $N$. We compare $S_{Q-C}$ to the entropy of the initial quartz phase, $S_{Q}$, with the same number of atoms $N$. $S_{Q-C}$ is larger than $S_{Q}$ because the entropy of the cristobalite phase is larger as discussed earlier. Hence,
$S_{Q-C}(T<T_{c})>S_{Q}(T<T_{c})$ (10)
The increased phase space and entropy in the high-temperature phase become “frozen-in” in the metastable state below $T_{c}$.
The quartz-cristobalite example is readily generalised to many systems where the combined entropy increases due to phase transitions. Figure 2 illustrates this point.
Figure 2: Quartz on the left transforms to cristobalite at high temperature as a result of a phase transition. In a set of solids, phase transitions result in larger phase space and higher entropy. Shown minerals are for illustration only.
3.2.2 Observability of phase transitions. Entropy without a phase transition
@@
We now observe that if we consider different values of FPCs, the phase transition would not operate. The upper bound on the maximal phonon frequency $\omega_{m}$ (“cage rattling” frequency) can be written in terms of fundamental constants Trachenko et al. . This bound follows from combining Eq. (4) and the ratio of $\hbar\omega_{m}$ and cohesive energy $E$ Trachenko and Brazhkin (2020)
@@
where $m$ is the atom mass. Setting $E=E_{\rm R}$ gives
@@
$\omega_{m}$ has the upper bound, $\omega_{m}^{u}$, for $m=m_{p}$ as Trachenko et al.
@@
Lets consider what would happen to a phase transition were FPCs to take different values and in particular consider variations of FPCs which increase $\omega_{m}$ and $\omega_{m}^{u}$ in Eqs. (12)-(13). Note that this question is well-posed because it is possible to vary $\omega_{m}$ and $\omega_{m}^{u}$ by varying $\hbar$, $m_{e}$, $e$ and $m_{p}$ while keeping $\beta$ and $\alpha$ in Eqs. (1) and (2) unchanged and hence maintain all preceding essential processes (e.g. those involved in stellar formation and evolution, synthesis of heavy elements, stability of ordered structures and so on) intact Barrow (2003); Adams (2019).
Large $\omega_{m}$ implies large $\omega_{i}$ and $\omega_{\rm D}$ (recall that $\omega_{m}$ is close to $\omega_{\rm D}$) and $T\ll\hbar\omega_{\rm D}$ in the temperature range considered. This, in turn, implies that the high-temperature classical result (5) does not apply because the system is now in the quantum regime. In this regime, instead of Eq. (5), the entropy is Landau and Lifshitz (1970)
@@
$S_{s}$ is close to 0 when $T\ll\hbar\omega_{\rm D}$. Then the second term in the free energy $F=E-TS$ is close to 0 too, and $F$ now does not have a crossover corresponding to a phase transition. The phase transition becomes inoperative.
The observability of phase transitions can also be discussed on the basis of $T_{c}$. In phase transition theories, including the Ising model and the anharmonic phonon theory (equivalent to Landau theory), the order of magnitude of $T_{c}$ is set by the energy of interaction between particles Dove (1997). This energy is on the order of the phonon energy $\hbar\omega_{\rm D}$ or $\hbar\omega_{m}$. Using Eq. (12), this gives
$T_{c}\approx\hbar{\omega_{m}}=\frac{1}{32\pi^{2}\epsilon_{0}^{2}}\frac{m_{e}e^{4}}{\hbar^{2}}\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ (15)
To estimate $T_{c}$, we write $m=Am_{p}$, where $A$ is the atomic mass. Using Eqs. (12) and (13), $T_{c}$ in Eq. (15) is
$T_{c}=\frac{\hbar\omega_{m}^{u}}{A^{\frac{1}{2}}}$ (16)
$A^{\frac{1}{2}}$ varies across the periodic table in the range of about 1 to 15, with an average value of 8. Using this value and $\omega_{m}^{u}=3680$ K in Eq. (13), Eq. (16) gives $T_{c}\approx$ 500 K. This is close to $T_{c}$ calculated from typical parameters of the anharmonic phonon model Dove (1993) and is of the right order of magnitude of $T_{c}$ in the range $10^{2}-10^{3}$ K typically observed.
We see that the order of magnitude of $T_{c}$ (15) is set by FPCs. Were different values of FPCs to give $T_{c}$ on the order of, for example, $10^{5}$ K or higher, phase transitions would be unobservable at our environmental conditions. This is consistent with our earlier argument involving small entropy in Eq. (14) due to large $\omega_{\rm D}$.
From this and previous sections, we conclude that the current values of FPCs, by enabling the operation of phase transitions which give rise to new states, increase the phase space and entropy.
3.2.3 Summary
The discussion in Section 3.2 gives us several insights. First, FPCs govern key effects in condensed matter physics and define the regime in which the system is, classical or quantum. Second, FPCs affect the observability and operation of entire effects such as phase transitions. Third, FPCs, by enabling phase transitions, increase the phase space and entropy.
@@ 3.4 First-order phase transitions: solids, liquids, gases
FPCs resulting in large cohesive energies and melting temperatures imply that matter would exist in the solid state only at given environmental conditions. Melting and boiling phase transitions would be unobservable. Conversely, FPCs resulting in small cohesive energies and hence low melting and boiling temperatures imply that matter would exist in the gas state only.
@@
Similarly to what we saw in Section 3.2 for the second-order transitions, the first-order phase transitions and the emerging high-temperature gas phase increase the entropy in comparison to what the entropy of the solid would have been in the absence of phase transitions (for first-order transitions, the entropy increase has a contribution from the latent heat). Similarly to the second-order transitions, the first-order transitions are enabled by FPCs as discussed earlier in this section.
Differently from solid and gas entropies we considered, discussing liquid entropy is more challenging theoretically due to the fundamental no-small parameter problem of the liquid theory (see Ref. Trachenko (2024) for a quick review and Refs. Trachenko and Brazhkin (2016); Trachenko (2023c) for a more detailed discussion). Common models such as the hard-sphere and Van der Waals models and their extensions give the constant-volume specific heat $c_{v}=\frac{3}{2}$ and hence describe the ideal gas thermodynamically rather than real liquids where $c_{v}=3$ close to melting Trachenko and Brazhkin (2016); Trachenko (2023c). However, the energy, the primary property in statistical physics Landau and Lifshitz (1970), can be calculated in liquids in the closed form using the main premise of statistical physics: the system’s energy is given by the energies of its excitations Lifshitz and Pitaevskii (2006). In liquids, these excitations are gapped phonons and particle jumps Trachenko and Brazhkin (2016); Trachenko (2023c); Proctor (2020, 2021); Chen (2022); de With (2024). The entropy can then be found by integrating $\frac{c_{v}}{T}$, where $c_{v}$ is calculated from the energy. Nevertheless, liquids are a condensed state of matter, and our discussion of how the interplay of cohesive energies set by FPCs and environmental parameters gives rise to different states of matter applies to liquids too.